Let V = Mn(C) and U be the subspace of all Hermitian matrices over C. Find a basis and

dimension of U.


1
Expert's answer
2021-01-15T15:03:17-0500

Hermitian matrix is a matrix A, which has components aij, that satisfy the condition "a_{ij} = \bar{a}_{ji}". This condition is invariant under multiplication by real numbers and not invariant under multiplication by complex numbers.

Therefore, U is not a complex subspace of Mn(C), but a real subspace only.

The real basis of U may be composed of the matrices of three types:

1) Dk, (k =1, 2, ..., n) which has zeros in all the cells besides the intersection of k-th column and k-th row, where it has 1. There are n matrices of this type.

2) Ekl, (k,l = 1, 2, ..., n, k>l) which has zeros in all the cells besides the intersection of k-th column and l-th row and the intersection of l-th column and k-th row, where it has 1. There are n(n-1)/2 matrices of this type.

3) Ikl, (k,l = 1, 2, ..., n, k>l) which has zeros in all the cells besides the intersection of k-th column, where it has i, and l-th row and the intersection of l-th column and k-th row, where it has -i. There are n(n-1)/2 matrices of this type.

The total amount of the matrices in this basis is n + n(n-1)/2 +n(n-1)/2 = n2.


Answer. "\dim_{\mathbb{R}} U = n^2"